# Faces of Fairness: Examining Bias in Facial Expression Recognition Datasets and Models

### v1

@@ 3.1 Data Analysis
To evaluate the intrinsic biases of each dataset, we designed an experiment where a model is trained to recognize the datasets. In this experiment, each sample was labeled with its corresponding dataset name, for both the training and validation sets. The model was then trained to predict the dataset of each image. High accuracy in this experiment indicates higher intrinsic biases, such as those related to illumination, background noise, camera resolution, gesture, eye gaze, head pose, hair color, and any other hidden bias in the datasets. In other words, a stronger correlation between any pair of datasets reflects a lower intrinsic bias in those datasets.
@@
We conducted an additional experiment to evaluate intrinsic biases by implementing a leave-one-dataset-out approach. In this experiment, we trained a deep model for FER using all except one dataset, which was held out for test. By assessing the model’s performance on the excluded dataset, we aimed to identify potential biases within the dataset itself while also measuring the model’s ability to generalize across different data sources. In this experiment, a lower accuracy suggests bias in the datasets. To ensure a fair comparison, we trained the model on the training sets of all datasets except one, validated its accuracy using their respective validation sets, and tested the bias using the validation set of the excluded (left-out) dataset. If the results of the validation and test experiments show similar accuracies, the model is considered unbiased. If there is a discrepancy, it indicates that the dataset suffers from bias.
To sum up, we extracted statistical data from each dataset and designed two model-based experiments—the dataset recognition model and the leave-one-dataset-out FER model—to assess both intrinsic and extrinsic biases within the datasets.
@@ 3.2 Model Analysis
A variety of general metrics have been proposed to evaluate fairness in machine learning models. Verma et al. [65] introduced three categories of fairness parameters: statistical measures, similarity-based measures, and causal reasoning. Additionally, Mittal et al. [66] presented metrics, such as parity-based metrics, score-based metrics, and facial-analysis-specific metrics.
@@
Recent studies have explored various bias and fairness metrics [20, 9, 69]. While some researchers introduced their own metrics, most studies have concentrated on widely recognized metrics such as Equalized Odds [44], Equal Opportunity [44], Demographic Parity [45, 46], Treatment Equality [47]. In this study, we focus on these four metrics and evaluate the bias and fairness of the models. Before delving deeper into these metrics, it is essential to explain the concepts of protected and unprotected groups. Protected groups are privileged against discrimination, while unprotected groups may face exclusion or bias.
Equalized Odds evaluates fairness by examining the positive predicted samples. Specifically, for both protected and unprotected groups, the model should exhibit similar ratios of True Positives (TP) and False Positives (FP). In other words, the likelihood of correctly identifying a positive case or making a False Positive prediction should not depend on the group membership.
@@
where $y\in\{0,1\}$. Equation 8 states that for an attribute $A$, a fair model will produce equal ratios of True Positives (TPs) and False Positives (FPs) for both the protected group ($A=0$) and unprotected group ($A=1$). For example, consider gender as the attribute, where $A=female$. A fair model should yield identical TP and FP ratios for females ($A=0$) and males ($A=1$). This ensures the model’s predictions are not disproportionately influenced by gender, satisfying Equalized Odds.
@@
Equal Opportunity focuses exclusively on the True Positive samples. Specifically, for a protected and an unprotected group, the probability of correctly predicting a positive outcome ($Y=1$) should be the same. The key difference between this metric and Equalized Odds is that, while Equalized Odds consider both True Positives ($Y=1$) and False Positives ($Y=0$), Equal Opportunity is solely concerned with the True Positive ($Y=1$).
Eq. 9 evaluates the equity in the ratio of TP between the protected ($A=0$) and unprotected ($A=1$) groups.
@@
Demographic Parity, also referred to as statistical parity, examines the independence of predictions from protected attributes. In essence, it evaluates the distribution of positive predictions, irrespective of whether they are TP or FP. This measure expects a uniform distribution of predictions across different groups. Eq. 10 provides a formal definition of Demographic Parity.
For instance, when considering the attribute race, the number of samples predicted as ${\hat{Y}}=1$ should be balanced across all racial groups, such as Black, White, and Asian.
@@
Treatment Equality examines the distribution of errors across different demographic groups. Specifically, it ensures that the ratio of False Negatives (FNs) to False Positives (FPs) remains consistent between protected and unprotected groups (see Eq. 11).
For example, the ratio of FN to FP should be consistent across all age groups, such as children, young adults, mature adults, and older individuals. Interestingly, some studies choose to evaluate this relationship by calculating the ratio of FPs to FNs instead.
@@
To obtain an evaluation score for fairness (bias score) in each model, we applied the four bias metrics to each expression label. Next, we selected the maximum bias across all expressions as the bias score for that metric (see Eq.12). To assess the bias of each attribute, we averaged the bias scores across the different metrics using Eq.13. Finally, the overall bias of each model was calculated by averaging the bias scores of the attributes (Eq. 14).
$Bias_{att}(met)=Max(Bias_{met,att}(exp))~{}~{}\forall exp\in EXP$, (12)
$Bias(att)=\frac{1}{K}\sum_{k=1}^{K}Bias_{att}(met)~{}~{}\forall met\in MET$, (13)
$Bias=\frac{1}{N}\sum_{n=1}^{N}Bias(att)~{}~{}\forall att\in ATT$. (14)
In Eq. 12, $EXP$ represents all seven basic expressions (Neutral, Happy, Sad, Surprise, Fear, Disgust, and Anger) [70]. In Eq. 13, $K$ denotes the number of bias metrics, while $MET$ refers to the individual metrics (Equalized Odds, Equal Opportunity, Demographic Parity, and Treatment Equality). In Eq. 14, $N$ represents the number of attributes, and $ATT$ refers to all attributes.
@@
To sum up, we employed four well-established bias metrics to evaluate key attributes in facial expression analysis, namely age, gender, and race. Our experiments were conducted on six popular models to assess their fairness. The detailed discussion of the models’ bias will follow in Section 4.
@@ 4 Experimental Results
This section explores bias across four well-known FER datasets: AffectNet [35], Fer2013 [37], RAF-DB [36], and ExpW [5]. Additionally, it examines the fairness in six distinct generations of deep models, including three CNN models: MobileNet [41], ResNet [42], XceptionNet [43], as well as three transformer-based models: ViT [32], CLIP [33], and GPT-4o-mini [34].
@@ 4.1 Bias in Datasets
Many studies suggest that the primary sources of bias in machine learning models lie in the data itself [10, 9, 11]. Consequently, to investigate bias in FER, it is essential to first examine the datasets. While numerous facial attributes can contribute to bias (see [60]), and a variety of biases can be present within the datasets (see [22, 9]), our focus is on the three most prominent facial attributes. In the following sections, we provide an overview of the datasets used in our study and detail the preprocessing steps we applied. Additionally, we explore the data distribution, attribute correlations, and model-based analysis.
@@ 4.1.1 Datasets
TABLE I: Diversity scores of the datasets is calculated using metrics: richness, evenness, and dominance. Scores are computed for each attribute, including expression, age, gender, and race. A weighted average score (with a richness coefficient of 0.25) is calculated for each attribute and dataset. The diversity score is the mean of the average scores. Notably, the data range from 0 to 100, with higher scores indicating greater diversity.
@@
Exp Age Gen Race Exp Age Gen Race Exp Age Gen Race Exp Age Gen Race
Richness 5.3 5.3 5.3 5.3 12.4 12.4 12.4 12.4 30.9 30.9 30.9 30.9 100 100 100 100
Evenness 84.2 52.9 96.4 69.7 93.0 54.5 95.7 58.0 77.1 54.4 87.0 76.6 74.0 52.8 97.3 70.8
Dominance 76.1 47.8 47.5 60.2 82.6 51.0 47.0 49.9 72.3 50.2 41.2 65.3 69.1 48.0 48.2 60.5
Average 53.9 34.0 48.4 43.7 59.6 36.2 48.6 37.0 52.4 37.4 45.3 49.9 56.0 41.9 56.8 52.1
Diversity 45 45.4 46.3 51.7
AffectNet is the largest in-the-wild FER dataset, comprising nearly 1 million images, with half of them manually annotated. It includes eight expressions (seven basic expressions plus Contempt), continuous labels (valence-arousal), and additional metadata. Fer2013 contains approximately 36K images, divided into 29K training samples and 3.5K validation samples. However, its small image size (40$\times$40 pixels) and grayscale color representation pose limitations. RAF-DB provides both single-label and compound-label subsets, with around 30K facial images and facial landmark annotations. The compound-labeling approach in RAF-DB results in 19 expressions, including the seven basic expressions and 12 compound expressions such as Happily-Surprise and Sadly-Angry. Finally, ExpW is a large-scale FER dataset, featuring over 90K manually annotated facial images. We utilize these datasets in our study to train and evaluate our models.
Figure 3: The data distribution across different datasets shows several trends: a) Happy and Neutral dominate the datasets, while Fear and Disgust are underrepresented. Among all datasets, Fer2013 exhibits the most balanced expression distribution, b) A noticeable bias is observed in the age groups, where [16$\sim$32] and [33$\sim$53] being more frequent, while [0$\sim$15] and [Over 54] have significantly fewer samples, c) Across all the datasets, there are more Man samples than Womans. This gender imbalance is most pronounced in the ExpW dataset and least evident in AffectNet, d) Regarding race, White samples are the most represented group, while Indians are the least represented. Data distribution for Black, Latinx, and Middle-Eastern races is more even.
@@ 4.1.2 Data Preprocessing
We used four in-the-wild datasets, each with different image sizes. In the first step, we aligned the images by ensuring the same cropping area, image size, and color domain across all the datasets. To achieve a consistent cropping area, we extended the original crop region by 25 to 35 percent in each dimension for all images. We then used the DeepFace model [71, 72] to process the extended crop and align the facial images across all datasets. Next, we resized all cropped faces to $224\times 224$. While three datasets are in the RGB color domain, Fer2013 is grayscale. To address this, we replicated the grayscale layer three times to match the RGB format, ensuring uniformity across the data. Additionally, since the ExpW dataset lacks a validation set, we randomly selected 10 percent of its images to serve as the validation set.
4.1.3 Data Diversity and Distribution
After data processing and achieving a unified format, we used the DeepFace model [68] to extract facial attributes, including age, gender, and race. Ages were classified into four ranges: [0$\sim$15], [16$\sim$32], [33$\sim$53], and [Over 54]. DeepFace predicts six racial categories: White, Black, Asian, Indian, Latinx, and Middle Eastern. For gender classification, the model outputs either Woman or Man. Throughout the remainder of this paper, we will use these terms to refer to the attribute values.
In our initial experiment, we assessed the diversity scores of the datasets using the metrics described in Sec. 3.1, namely Richness, Evenness, and Dominance. By applying Eqs. 12, and 3, we derived the diversity scores summarized in Table I. This table presents the diversity scores for each dataset, where the average score is a weighted combination of the attributes. Notably, Richness was assigned a weight of 0.25 because it is a dataset-level metric rather than an attribute-specific one, resulting in the same score across all attributes. The overall diversity score was then calculated as the mean of these weighted averages. Our findings indicate that AffectNet is the most diverse dataset, whereas the other datasets scored below 50 on average out of 100.
Figure 4: The 2D correlation matrix illustrates the relationships between different attributes across all the datasets (in percent). The diagram visualizes the data distribution for each attribute value and highlights the contribution of each attribute to others. This heat map reveals biases toward Neutral and Happy. The age groups [16$\sim$32] and [33$\sim$53] are the most prominent, while Man appears the most frequent gender. In terms of race, White is overrepresented, whereas Indian is underrepresented.
In a subsequent experiment, we applied global-and-local-stereotypical metrics (Eq. 4) to evaluate data bias. Fig. 3-a depicts the expression distribution of datasets, highlighting that Neutral and Happy are the most prevalent, whereas Fear and Disgust are the least represented. Notably, Fer2013 and RAF-DB demonstrate a more balanced distribution of expressions compared to the pronounced imbalances observed in AffectNet and ExpW. Fig. 3-b reveals a pronounced bias in dataset across age groups, with the [16$\sim$32] and [33$\sim$53] ranges being significantly overrepresented. In contrast, samples from the [0$\sim$15] and [Over 54] age groups account for less than 5% of the total data in all datasets.
@@ 4.1.3 Data Diversity and Distribution
As shown in Fig. 3-c, the number of Man samples is approximately 50% more than Woman, though this gender bias is less pronounced compared to other attributes. Fig. 3-d, on the other hand, highlights a notable racial imbalance, with White samples constituting over 50% of the total across all datasets. The distribution of Asian, Black, Latinx, and Middle Eastern samples is more balanced, while Indian samples remain consistently underrepresented. In conclusion, an examination of the data distribution across four essential attributes—age, gender, race, and expression—uncovers notable imbalances that may introduce bias into the model. Addressing these disparities is vital to ensuring fairness in FER systems.
@@ 4.1.4 Attributes Correlation
Building upon the global-and-local-stereotypical analysis, we applied Eq. 5 to explore bias across the data samples. Additionally, examining the correlation between different attributes offers another meaningful approach for data analysis. For this purpose, we focused on expressions as the primary attribute and analyzed their correlation with other facial attributes of age, gender, and race.
Figure 4 illustrates that 27.7% of the data is labeled Neutral, where 18.1% belong to the [16$\sim$32] age group, and 9.3% fall within [33$\sim$53]. In contrast, less than 0.3% of Neutral label corresponds to the two remaining age groups. Regarding gender distribution, Neutral reveals a significant disparity between Man and Woman samples, with the number of Man being more than double that of Woman. For racial distribution within this expression, 15.5% of the samples are associated with White, while the total representation of the other five racial groups is below 13%.
The distribution of Happy exhibits a similar bias toward the age groups [16$\sim$32] and [33$\sim$53]. However, no significant gender disparity is observed between Man and Woman. In contrast, bias exists toward White group for this expression. Compared to Happy and Neutral, the bias across attributes is less pronounced for the other expressions, particularly across age groups and racial categories. Among the datasets, ExpW demonstrates the highest bias, while Fer2013 emerges as the least biased.
@@
The joint attribute correlation, calculated using Eqs. 6 and 7, offers deeper insights into bias in dataset. Fig. 5 reveals that the Happy and Neutral expressions, [16$\sim$32] and [33$\sim$53] age groups, Man, and White are the most correlated values across attributes. Specifically, Happy and Neutral account for 60.7% of the data, 98.6% of the samples fall within the [16$\sim$53] age range, 62.6% of the samples are labeled as Man, and White dominates with 59.3% of the racial distribution.

### v2

@@ 3.1 Data Analysis
We also evaluated non-demographic biases in each dataset. To do so, we designed an experiment where a model is trained to recognize the datasets. In this experiment, each image sample was labeled with its corresponding dataset name, for both the training and validation sets. Then, the model was trained to predict the dataset of each image. High accuracy in this experiment indicates higher non-demographic bias, such as those related to illumination, background noise, camera resolution, gesture, eye gaze, head pose, hair color, and any other hidden bias in the datasets.
@@
In addition, we conducted another experiment to evaluate bias in datasets. In this leave-one-dataset-out experiment, we trained a deep model for facial expression recognition, using all except one dataset, which was left out for test. By assessing the model’s performance on the excluded dataset, we aimed to identify potential biases within the dataset. In this experiment, lower accuracy suggests higher bias in the datasets. To ensure a fair comparison, we trained the model on the training sets of all datasets except one, validated its accuracy using their respective validation sets, and tested the bias using the validation set of the excluded (left-out) dataset. If the results of the validation and test experiments show similar accuracies, the model is considered unbiased. Any discrepancy indicates that the dataset suffers from bias. To sum up, we utilized mathematical data analysis and designed two model-based experiments— dataset recognition model and leave-one-dataset-out FER model—to assess bias within the datasets.
@@ 3.2 Model Analysis
A variety of general metrics have been proposed to evaluate fairness in machine learning models [33, 9, 54].Verma et al. [52] introduced three categories of fairness parameters: statistical measures, similarity-based measures, and causal reasoning. Additionally, Mittal et al. [53] introduced metrics, such as parity-based metrics, score-based metrics, and facial-analysis-specific metrics.
@@
While some researchers introduced their own metrics, most studies have concentrated on widely recognized metrics such as Equalized Odds [55], Equal Opportunity [55], Demographic Parity [56, 57], and Treatment Equality [58]. In this study, we focus on these four metrics to evaluate bias and fairness of models. To study bias in the models, we need to know the concept of protected and unprotected groups. Protected groups are privileged against discrimination, while unprotected groups may face exclusion or bias. In continue, we quickly review the four bias metrics we are going to utilize in this research.
@@
Equalized Odds evaluates fairness by examining the positive predicted samples. Eq. 8 states that for an attribute $A$, a fair model will produce equal ratios of True Positives (TPs) and False Positives (FPs) for both the protected group ($A\mathord{=}a$) and unprotected group ($A\mathord{=}b$):
$\begin{aligned} p(\hat{Y}=1|A=a,Y=y)=p(\hat{Y}=1|A=b,Y=y).\end{aligned}$
@@
Equal Opportunity focuses exclusively on TPs, where the probability of TPs ($Y\mathord{=}1$) for protected and an unprotected groups should be the same. The key difference between this metric and Equalized Odds is illustrated in Eq. 9. While Equalized Odds consider both TPs and FPs, Equal Opportunity concentrates on TPs:
$\begin{aligned} p(\hat{Y}=1|A=a,Y=1)=p(\hat{Y}=1|A=b,Y=1).\end{aligned}$
@@
Demographic Parity examines the independence of predictions from attributes. In essence, it evaluates the distribution of positive predictions, irrespective of whether they are TP or FP. As Eq. 10 shows, this metric expects a uniform distribution of predictions across different groups:
$\begin{aligned} p(\hat{Y}=1|A=a)=p(\hat{Y}=1|A=b).\end{aligned}$
@@
Treatment Equality examines the distribution of errors across different demographic groups (see Eq. 11). Specifically, it ensures that the ratio of False Negatives (FNs) to False Positives (FPs) remains consistent between protected and unprotected groups:
$\begin{aligned} \frac{p({\hat{Y}}=0|A=a,Y=1)}{p({\hat{Y}}=1|A=a,Y=0)}=\frac{p({\hat{Y}}=0|A=b,Y=1)}{p({\hat{Y}}=1|A=b,Y=0)}.\end{aligned}$
@@
To obtain bias score of each model, for each demographic attribute, we applied all four bias metrics to each expression label. Next, we selected the maximum bias across all expressions as the bias score for that metric (see Eq.12). To assess the bias of each attribute, we averaged the bias scores across the different metrics using Eq.13. Finally, the overall bias of each model was calculated by averaging the bias score of each attribute (Eq. 14):
$\begin{aligned} Bias_{att}(met)=Max(Bias_{met,att}(y))~~~~\forall~y\in Y,\end{aligned}$
$\begin{aligned} Bias(att)=\frac{1}{4}\sum Bias_{att}(met)~~~~\forall~met\in MET,\end{aligned}$
$\begin{aligned} Bias=\frac{1}{3}\sum Bias(att)~~~~\forall~att\in ATT.\end{aligned}$
@@
To sum up, we employed four well-established bias metrics to evaluate fairness of eight FER models over three demographic attributes, age, gender, and race. The detailed discussion of the models’ bias will follow in Section 4.
@@ 4 Experimental Results
This section explores bias across four well-known FER datasets: AffectNet [25], Fer2013 [27], RAF-DB [26], and ExpW [5]. Additionally, it examines fairness in eight generations of deep models, including three CNN models: MobileNet [28], ResNet [29], XceptionNet [30], as well as three transformer-based models: ViT [22], CLIP [23], and GPT-4o-mini [24], plus two state-of-the-art FER-specialized models: POSTER [31] and CEPrompt [32].
@@ 4.1 Bias in Datasets
Many studies suggest that the primary sources of bias in machine learning models lie in the data itself [10, 9, 11]. Consequently, to investigate bias in FER, it is essential to first examine the datasets. While numerous facial attributes can contribute to bias [59], and a variety of biases can be present within the datasets [41, 9], our focus is on the three most prominent facial attributes. In the following sections, we provide an overview of the datasets used in our study and detail the preprocessing steps we applied. Additionally, we explore the data distribution, attribute correlations, and model-based analysis.
@@ 4.1.1 Datasets
TABLE II: Bias metrics for Age, Gender, Race, and their mean across datasets (in percent), include Wasserstein Distance (WD), Jensen-Shannon Divergence (JSD), Conditional-Entropy Bias Index (CEBI), Sparsity Index (SI), Normalized Shannon Entropy (NSE), Normalized Label Skewness (NLS), and Group Normalized Mutual Information (GNMI) metrics.
@@
Age Gender Race Mean Age Gender Race Mean Age Gender Race Mean Age Gender Race Mean
WD 5.3 3.1 3.1 3.8 4.9 3.6 3.6 4.0 5.0 6.0 2.7 4.6 5.3 2.7 3.0 3.6
JSD 2.2 0.4 1.0 1.2 2.2 0.5 1.3 1.3 2.2 1.1 0.7 1.3 2.2 0.4 0.9 1.2
CEBI 82.2 63.7 90.5 78.8 83.1 64.8 89.4 79.1 82.0 61.5 91.0 78.2 81.6 62.5 90.3 78.2
SI 13.0 9.5 14.8 12.4 14.1 11.2 15.0 13.4 12.6 7.0 14.8 11.5 11.9 7.4 14.2 11.2
NSE 5.2 10.3 3.5 6.3 8.3 16.3 5.5 10.0 2.8 5.2 1.8 3.3 1.6 5.4 1.1 2.7
NLS 54.0 51.4 51.8 52.4 30.0 49.9 32.2 37.4 47.6 50.1 46.8 48.2 58.5 54.2 51.3 54.7
GNMI 61.9 29.3 76.0 55.8 54.7 26.1 70.2 50.4 64.4 31.7 83.7 59.9 67.1 33.3 84.3 61.6
Bias 30.1 27.9 29.6 30.5
AffectNet is the largest in-the-wild FER dataset, comprising nearly 1 million images, with half of them manually annotated. It includes eight expressions (seven basic expressions plus Contempt), continuous labels (valence-arousal), and additional metadata. Fer2013 contains approximately 36K images, divided into 29K training samples and 3.5K validation samples. However, its small image size ($\times40$ pixels) and grayscale color representation pose limitations. RAF-DB provides both single-label and compound-label subsets, with around 30K facial images and facial landmark annotations. The compound-labeling approach in RAF-DB results in 19 expressions, including the seven basic expressions and 12 compound expressions such as Happily-Surprise and Sadly-Angry. Finally, ExpW is a large-scale FER dataset, featuring over 90K manually annotated facial images. We utilize these datasets in our study to train and evaluate our models.
[…]
@@ 4.1.2 Data Preprocessing
To have a unique form of data, in the first step, we aligned images by ensuring the same cropping area, image size, and color domain across all the datasets. We extended the original crop region by 25 to 35 percent in each dimension for all images, then used the DeepFace model [61, 62] to process the extended crop and align the facial images. Next, we resized all the cropped faces to $224\times 224$. Since Fer2013 is grayscale, we replicated the grayscale layer three times to match the RGB format, ensuring uniformity across the data. Additionally, ExpW lacks a validation set, therefore, we randomly selected 10 percent of its images to serve as the validation set.
4.1.3 Metric-Based Dataset Analysis
[…]
In the initial experiment, we assessed dataset bias using the metrics described in Sec. 3.1. By applying Eqs. 1 to 7, we derived the bias scores summarized in Table II. This table presents the scores for each metric and demographic attribute, where the final bias score of each dataset is the average of different scores.
The bias analysis of the four datasets reveals notable demographic disparities, with overall bias scores ranging from 27.9% (Fer2013) to 30.5% (AffectNet). None of the datasets are fully balanced across Age, Gender, and Race, suggesting that models trained on these datasets are likely to inherit these imbalances. Among the attributes, Race consistently exhibits the highest bias values across most metrics, indicating that racial representation is more uneven than Age and Gender. Gender shows comparatively lower bias, although metrics like CEBI and NLS indicate persistent disparities, while Age-related bias is moderate but consistent. These findings highlight that demographic representation is far from uniform, and differences in attribute distribution could affect fairness in model predictions.
Examining individual metrics provides deeper insight into the nature of these biases. Metrics such as CEBI and GNMI report high values across all datasets, indicating strong conditional dependencies and unequal mutual information between demographic attributes and labels. SI values reveal the uneven distribution of samples across different attribute-label combinations, while NSE and NLS reflect deviations from uniform label distributions, with Fer2013 showing higher unevenness in Age and Gender, and AffectNet showing pronounced skew in Age. WD and JSD are relatively lower, suggesting that while overall distributional differences exist, they are less pronounced than the conditional dependencies captured by CEBI and GNMI. Overall, these results underscore the importance of considering both dataset composition and metric-specific insights when addressing bias in facial expression recognition.
[…]
4.1.4 Attributes Distribution
[…]
@@ 4.1.4 Attributes Distribution
As shown in Fig. 3-c, the number of Man samples is approximately 50% more than Woman, though this gender bias is less pronounced compared to other attributes. Fig. 3-d, on the other hand, highlights a notable racial imbalance, with White samples constituting over 50% of the total across all datasets. The distribution of Asian, Black, Latinx, and Middle Eastern samples is more balanced, while Indian samples remain consistently underrepresented. In conclusion, an examination of the data distribution across four essential attributes—age, gender, race, and expression—uncovers notable bias in the datasets, where addressing them is vital to ensure fairness in FER models.
@@ 4.1.5 Attributes Correlation
Building on Micro-Level Correlation Bias, we studied the joint probability distribution of data. Examining correlations between joint attributes provides a meaningful approach for data analysis. We kept expression as the primary attribute and analyzed its correlation with other facial attributes: age, gender, and race.
[…]
@@
[…]